# ReaGAN: Node-as-Agent-Reasoning Graph Agentic Network

### v2

@@ Abstract
Graph Neural Networks (GNNs) have achieved remarkable success in graph-based learning by propagating information among neighbor nodes by predefined aggregation mechanisms. However, such fixed schemes often suffer from two key limitations. On the one hand, it cannot handle the imbalance in node informativeness—some nodes are rich in information, while others remain sparse. On the other hand, predefined message passing primarily leverages local structural similarity while ignoring global semantic relationships across the graph, limiting the model’s ability to capture distant but relevant information. To address these limitations, we propose Retrieval-augmented Graph Agentic Network (thus named ReaGAN), an agent-based framework that addresses these limitations by empowering each node with autonomous, individual node-level decision-making. Each node is treated as an agent that independently plans its next action based on its internal memory, enabling node-level planning and adaptive message propagation. In addition, retrieval-augmented generation (RAG) is used to enable nodes to retrieve semantically relevant content and build the global relationship among the nodes in the graph. Extensive experiments demonstrate that ReaGAN achieves competitive performance under few-shot in-context settings, using only a frozen LLM backbone without fine-tuning. These results highlight the potential of agentic planning and integrated local-global retrieval for advancing graph machine learning.
1 Introduction
@@
Graph Machine Learning (GML) has achieved remarkable success over the past years, with Graph Neural Networks (GNNs) such as GCN (Kipf and Welling, 2017), GAT (Veličković et al., 2018), and GraphSAGE (Hamilton et al., 2017) becoming the de facto standards for representation learning over graph-structured data. These models operate through a static, globally synchronized message-passing framework, where in each layer, every node aggregates information from its neighbors using a predefined aggregation-update rule parameterized by shared weights across the graph.
@@
Figure 1: Message passing in traditional pre-defined way vs. ReaGAN’s method. (a) All nodes only do Local Aggregation. (b) Nodes aggregate in different ways. A’s is local and global; B’s is local only; C’s is global only; D does no operation.
@@
While effective in many scenarios, this paradigm suffers from a fundamental limitation: it treats all nodes uniformly, regardless of their varying local context or inherent semantics (Finkelshtein et al., 2023; Martinkus et al., 2023). The aggregation-update process is entirely homogeneous, where each node follows the same procedure, every layer applies the same function, and no node has the capacity for individualized decision-making. However, graphs often contain a small subset of nodes that are rich in semantic content or structurally well-positioned, while others are sparsely connected, noisy, or contextually ambiguous. Thus, applying identical message-passing rules to all nodes in such settings is problematic: highly informative nodes may be overwhelmed by irrelevant inputs, while uninformative nodes fail to gather the support they need. Worse still, this homogeneous propagation can amplify noise and redundancy, thereby degrading overall representation quality. This observation motivates Challenge 1: Node-Level AutonomyCan we endow each node with the capacity to autonomously plan its own message-passing behavior based on its internal state and local context, rather than relying on globally shared rules?
@@
In addition, most existing GNNs operate primarily based on local structural similarity, implicitly assuming that neighboring nodes share informative and task-relevant features. However, this assumption often breaks down in real-world graphs where semantically related nodes may be structurally distant, which is a pattern especially common in open-domain or heterogeneous networks (Zhu et al., 2024; Li et al., 2024). As a result, traditional GNNs struggle to capture global semantic dependencies, limiting their ability to perform long-range reasoning or generalize beyond local structure. This limitation is especially pronounced in sparsely connected regions, where the local neighborhoods offer limited predictive value Li et al. (2024). In such settings, retrieving semantically aligned but distant nodes becomes essential for enriching context and improving node representations. This raises Challenge 2: Local-Global ComplementarityCan we combine local-structure neighbors with global-semantic neighbors to enable a more comprehensive context-aware message passing?
@@
To address the two key challenges, we propose a new perspective on graph learning by rethinking the fundamental computational unit: we treat each node as an autonomous agent. Unlike GNNs that passively aggregate messages through static rules, each node acts as an agent who actively determines its action at each layer based on internal states and contextual signals. Inspired by general agent-based systems (Russell and Norvig, 2010; Wang et al., 2023; Xu et al., 2025; Sun et al., ; Gao and Zhang, 2024), our agentic framework endows each node with four core components: Planning, which decides the next operations based on its current objective and context; Actions, which executes local-global aggregation strategies to interact with neighbors (Yao et al., 2022; Shinn et al., 2023); Memory, which stores the node’s cumulative textual features along with historical interaction traces; and Tool Use, which calls external functions like retrieval-augmented generation. In this work, we unify these components into the Retrieval-augmented Graph Agentic Network (ReaGAN), which enables each node to make individualized, adaptive, and semantically informed decisions. This self-decision design directly addresses Challenge 1 by granting each node autonomy in its message passing, while the local-global aggregation strategy addresses Challenge 2 as the following paragraph shows.
@@
Beyond Local Aggregation which enables nodes to exchange messages with directly connected neighbors, we integrate retrieval-augmented generation (RAG) (Lewis et al., 2020; Su et al., 2025) as an external tool that empowers nodes to access semantically relevant but structurally distant information, where the entire graph is viewed as a searchable database. Each node then integrates both local structural signals and globally retrieved semantic content into its individual memory, forming a richer contextual representation for subsequent planning and interaction. Equipped with these capabilities, each node engages in an agentic workflow. Based on its current state and memory, the node prompts a frozen LLM to generate a context-aware plan, selects and executes the appropriate actions, such as local or global aggregation, optionally invokes external tools like RAG, and subsequently updates its internal memory with the acquired information.
@@
We pinpoint the key differences between traditional message passing and our agentic workflow. As shown in Figure  1(a), standard GNNs enforce a uniform propagation rule, where all nodes solely communicate only with immediate neighbors. In contrast, Figure  1(b) illustrates ReaGAN’s mechanism, in which each node autonomously selects its own actions and possibly reach its remote yet semantical neighbors. For example, node A performs both local and global aggregation; node B conducts only local aggregation, resembling classical GNN behavior; node C executes only global aggregation; and node D chooses to remain inactive. This demonstrates how ReaGAN supports diverse and asynchronous strategies at the node level, offering a level of node-level flexibility.
@@
$\bullet$ We introduce ReaGAN, an agentic graph learning framework that models each node as an autonomous agent equipped with planning, memory, action, and tool-use capabilities, moving beyond static, rule-based message passing in traditional GNNs.
$\bullet$ We introduce a hybrid aggregation mechanism that integrates local structural and global semantic information via retrieval-augmented generation, allowing nodes to dynamically access semantically relevant but structurally distant context.
$\bullet$ We conduct extensive experiments to demonstrate that ReaGAN achieves strong performance over existing traditional methods, even using a frozen LLM without fine-tuning.
2 Method
Node As Agent
In ReaGAN, each node is treated as an autonomous agent capable of perceiving its own state and neighborhood context, planning its next steps, executing context-aware actions, utilizing external tools, and updating its internal memory. This perspective departs from traditional synchronized message passing, instead enabling fully individualized, asynchronous, and adaptive behavior. Moreover, each node can not only aggregate information from local structural neighbors, but also retrieve semantically relevant signals from distant nodes across the graph. By doing so, ReaGAN naturally supports both node-level personalization and autonomy (addressing Challenge 1) and joint local-global information integration (addressing Challenge 2).
Agentic Formulation
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be an attributed graph, where each node $v\in\mathcal{V}$ is associated with a text feature $t_{v}$ and an optional label $y_{v}\in\mathcal{Y}$. The goal is to predict labels $\hat{y}_{v}$ for all unlabeled nodes. Each node is treated as an agent equipped with memory $\mathcal{M}_{v}$, interacting with a frozen large language model (LLM) through a multi-layered reasoning loop. At each layer, the node constructs a prompt based on its memory, queries the LLM for an action plan, and updates its memory accordingly. After $L$ layers, the node predicts a label by querying the LLM with a prompt derived from its final memory state.
In summary, each node follows a layer-wise cycle of perception, planning, action execution, and memory update, independently deciding whether to gather local/global information, make a prediction, or take no action. This fully agentic workflow is detailed in Algorithm 1.
@@ 2 Method
16:   else if $a=\texttt{Prediction}$ then
17:    $p_{v}^{\texttt{pred}}\leftarrow\texttt{Prompt}_{\texttt{predict}}(\mathcal{M}_{v}^{(l-1)})$
18:    $\hat{y}_{v}\leftarrow\texttt{LLM}(p_{v}^{\texttt{pred}})$
19:   else if $a=\texttt{NoOp}$ then
20:    $\mathcal{M}_{v}^{(l)}\leftarrow\mathcal{M}_{v}^{(l-1)}$ {No change to memory}
21:   end if
22:  end for
23:end for
24:return  Predicted label $\hat{y}_{v}$ (if generated)
2.1 Planning: Prompting Frozen LLM
In ReaGAN, each node is equipped with a local planner that determines how to act at each layer. Rather than relying on a globally shared aggregation rule, we leverage a frozen large language model (LLM) to enable in-context planning, where decisions are conditioned on the node’s own memory. At each layer $l$, node $v$ constructs a structured prompt based on its internal memory $\mathcal{M}_{v}^{(l)}$. The prompt is then sent to a frozen LLM (e.g., LLaMA (Touvron et al., 2023), Qwen (Zhou et al., 2024), or DeepSeek (Team, 2024)), which returns an action plan:
$a_{v}^{(l)}=\texttt{LLM}(\texttt{Prompt}_{\texttt{planning}}(\mathcal{M}_{v}^{(l)}))$
while the action(s) in $a_{v}^{(l)}$ are selected from a discrete space and described in detail in Section 2.2. The node then parses and executes the plan, updating its memory with newly acquired information. This process is repeated for $L$ layers. At the final layer, the node constructs a prediction-specific prompt and queries the LLM to produce a label:
@@ 2.1 Planning: Prompting Frozen LLM
which ends up the overall planning loop illustrated in Figure 2. To sum up, the planning process allows each node to reason independently and asynchronously, determining what to do and when to act—without any global synchronization. It forms the core mechanism that enables Challenge 1 (Node-Level Autonomy) and Challenge 2 (Global Semantic Access) to be addressed in a unified and decentralized manner.
Figure 2: Overview of ReaGAN. Each node in ReaGAN is modeled as an agent equipped with four core modules: Memory, Planning, Tools, and Action. The node stores its original information and receives local and global information into its memory, constructs a natural language prompt, and queries a frozen LLM for an action plan. The returned plan may contain one or more actions (e.g., Local Aggregation, Global Aggregation), which are executed using available tools such as RAG. The resulting outputs are written back into memory and may produce a predicted label. This forms a closed loop of perception, planning, action, and memory refinement across layers.
2.2 Action: Node-Level Decision Space
@@
Both the aggregated summary and the selected labeled examples are written into the node’s memory, which is same as the local version. Through this process, Global Aggregation provides semantic enrichment and label grounding from structure-agnostic sources. It expands each node’s informational horizon beyond its local neighborhood—especially benefiting nodes in sparse or isolated regions. This mechanism directly addresses Challenge 2 (Global Semantic Access).
2.2.3 NoOp.
@@
This mechanism is essential for avoiding information over-collection, especially in cases where the memory already contains sufficient context. It helps prevents noise accumulation and supports pacing in multi-layer reasoning. By allowing nodes to wait or opt out of message passing altogether, NoOp reinforces ReaGAN’s core principle of self-decision and resource-aware adaptation (Challenge 1).
2.3 Memory: Tracking the Internal State
@@
where $\{m_{i}^{(l)}\}$ are the new entries produced by actions such as Local Aggregation, Global Aggregation, or retrieval-based example selection at layer $l$, which is equals to $\tilde{t}_{v}^{(l)}\ \cup\mathcal{E}_{v}^{(l)}$).
@@
Figure 3: Information flow from memory to prompt. Each node’s memory includes its original text feature, aggregated text feature from local and global neighbors, and selected labeled neighbor shots. During planning or prediction, these components are selectively injected into a natural language prompt, providing the LLM with (i) the input node’s raw identity, (ii) context-enhanced descriptions, and (iii) label-text pairs for few-shot learning. This design enables each agent to reason over multi-scale context and take personalized actions.
2.4 Tools: Global Semantic Augmentation

### v4

@@ Abstract.
Graph Neural Networks (GNNs) have achieved remarkable success in graph-based learning by propagating information among neighbor nodes through predefined aggregation mechanisms. However, such fixed schemes often suffer from two key limitations. On the one hand, it cannot handle the varying node informativeness—some nodes are rich in information, while others remain sparse. On the other hand, predefined message passing primarily leverages local structural similarity while ignoring global semantic relationships across the graph, limiting the model’s ability to capture distant but relevant information. To address these limitations, we propose the Retrieval-augmented Graph Agentic Network (ReaGAN), an agent-based framework that addresses these limitations by empowering each node with autonomous, individual node-level decision-making. Each node is treated as an agent that independently plans its next action based on its internal memory, enabling node-level planning and adaptive message propagation. Furthermore, we integrate retrieval-augmented generation (RAG) as a tool for agents to dynamically build global relationships by retrieving semantically relevant content from across the graph. Extensive experiments demonstrate that ReaGAN achieves competitive performance under few-shot in-context settings, using only a frozen LLM backbone without fine-tuning. These results highlight the potential of agentic planning and integrated local-global retrieval for advancing graph machine learning.
Graph Neural Networks, Large Language Models, Retrieval-Augmented Generation, Agent-based Graph Learning
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Artificial intelligenceccs: Information systems Information retrieval
1. Introduction
@@
Graph Machine Learning (GML) has achieved remarkable success over the past years, with Graph Neural Networks (GNNs) such as GCN (Kipf and Welling, 2017), GAT (Veličković et al., 2018), and GraphSAGE (Hamilton et al., 2017) becoming the de facto standards for representation learning over graph-structured data. These models operate through a static, globally synchronized message-passing framework, where in each layer, every node aggregates information from its neighbors using a predefined aggregation and update rule parameterized by shared weights across the graph.
@@
Figure 1. Message passing in traditional pre-defined way vs. ReaGAN’s method. (a) All nodes perform local aggregation only. (b) Nodes aggregate in different ways. A aggregates both locally and globally; B performs local-only aggregation; C performs global-only aggregation; D performs no operation.
@@
While effective in many scenarios, this paradigm suffers from a fundamental limitation: it treats all nodes uniformly, regardless of their varying local context or inherent semantics (Finkelshtein et al., 2023; Martinkus et al., 2023). The aggregation-update process is homogeneous: each node follows the same layer-wise rule, and no node has the capacity for individualized decision-making. However, graphs often contain a small subset of nodes that are rich in semantic content or structurally well-positioned, while others are sparsely connected, noisy, or contextually ambiguous. Thus, applying identical message-passing rules to all nodes in such settings is problematic: highly informative nodes may be overwhelmed by irrelevant inputs, while uninformative nodes fail to gather the support they need. Worse still, this homogeneous propagation can amplify noise and redundancy, thereby degrading overall representation quality. This observation motivates Challenge 1: Node-Level AutonomyCan we endow each node with the capacity to autonomously plan its own message-passing behavior based on its internal state and local context, rather than relying on globally shared rules?
@@
In addition, most existing GNNs operate primarily based on local structural similarity, implicitly assuming that neighboring nodes share informative and task-relevant features. However, this assumption often breaks down in real-world graphs where semantically related nodes may be structurally distant, which is a pattern especially common in open-domain or heterogeneous networks (Zhu et al., 2024; Li et al., 2024). As a result, traditional GNNs struggle to capture global semantic dependencies, limiting their ability to perform long-range reasoning or generalize beyond local structure. This limitation is especially pronounced in sparsely connected regions, where the local neighborhoods offer limited predictive value (Li et al., 2024). In such settings, retrieving semantically aligned but distant nodes becomes essential for enriching context and improving node representations. This raises Challenge 2: Local-Global ComplementarityCan we combine local-structure neighbors with global-semantic neighbors to enable a more comprehensive context-aware message passing?
@@
To address the two key challenges, we propose a new perspective on graph learning by rethinking the fundamental computational unit: we treat each node as an autonomous agent that decides its own aggregation actions. Unlike GNNs that passively aggregate messages through static rules, each node acts as an agent that actively determines its action at each layer based on internal states and contextual signals. Inspired by general agent-based systems (Russell and Norvig, 2010; Wang et al., 2023; Xu et al., 2025; Sun et al., [n. d.]; Gao and Zhang, 2024; Li et al., 2025a; Nie et al., 2024), our agentic framework endows each node with four core components: Planning, which decides the next operations based on its current objective and context by a frozen LLM; Actions, which execute local-global aggregation strategies to interact with neighbors (Yao et al., 2022; Shinn et al., 2023); Memory, which stores the node’s cumulative textual features along with selected neighbor features; and Tool-Use, which calls external functions like retrieval-augmented generation to enhance the feature representation. In this work, we unify these components into the Retrieval-augmented Graph Agentic Network (ReaGAN), which enables each node to make individualized, adaptive, and semantically informed decisions. This self-decision-making design directly addresses Challenge 1, while the local-global aggregation strategy addresses Challenge 2 as the following paragraph shows.
@@
Beyond Local Aggregation, which enables nodes to exchange messages with directly connected neighbors, we integrate retrieval-augmented generation (RAG) (Lewis et al., 2020; Su et al., 2025) as an external tool that empowers nodes to access semantically relevant but structurally distant information, where the entire graph is viewed as a searchable database. Each node then integrates both local structural signals and globally retrieved semantic content into its memory, forming a richer contextual representation for subsequent planning and interaction. Equipped with these capabilities, each node engages in an agentic workflow. Based on its current state and memory, the node prompts a frozen LLM to generate a context-aware plan, selects and executes the appropriate actions, such as local or global aggregation, optionally invokes external tools like RAG, and subsequently updates its internal memory with the acquired information.
@@
We pinpoint the key differences between traditional message passing and our agentic workflow. As shown in Figure  1(a), standard GNNs enforce a uniform propagation rule, where all nodes communicate only with immediate neighbors. In contrast, Figure  1(b) illustrates ReaGAN’s mechanism, in which each node autonomously selects its own actions and possibly access semantically similar but structurally distant neighbors. For example, node A performs both local and global aggregation; node B conducts only local aggregation, resembling classical GNN behavior; node C executes only global aggregation; and node D chooses to remain inactive. This demonstrates how ReaGAN supports diverse and asynchronous strategies at the node level, offering node-level flexibility.
Crucially, our entire agentic framework is powered by a frozen, off-the-shelf Large Language Model (LLM) that serves as the core reasoning engine for each node agent. This design choice is fundamental to our contribution and carries significant implications. By leveraging a pre-trained and frozen LLM, we completely bypass the need for costly, gradient-based training and task-specific fine-tuning, which are hallmarks of traditional GNNs. This approach positions ReaGAN not merely as a new architecture, but as a more resource-efficient and flexible “plug-and-play” paradigm for graph learning. It allows us to directly tap into the powerful, generalized reasoning capabilities of the LLM, enabling strong performance in few-shot scenarios without learning any trainable parameters for the graph task at hand.
@@
  • We introduce ReaGAN, an agentic graph learning framework that models each node as an autonomous agent equipped with planning, memory, action, and tool-use capabilities, thereby moving beyond the static, rule-based message passing inherent in traditional GNNs.
  • We introduce a hybrid aggregation mechanism that integrates local structural and global semantic information via retrieval-augmented generation, allowing nodes to dynamically access semantically relevant but structurally distant context.
  • We conduct extensive experiments to demonstrate that ReaGAN achieves competitive performance compared to existing traditional supervised methods, even when using a frozen LLM without fine-tuning.
2. Method
2.1. Node as Agent
In ReaGAN, each node is treated as an autonomous agent capable of perceiving its own state and neighborhood context, planning its next steps, executing context-aware actions, utilizing external tools, and updating its internal memory. This perspective departs from traditional synchronized message passing, instead enabling fully individualized, asynchronous, and adaptive behavior. Moreover, each node can aggregate from local structural neighbors, and retrieve from global semantic neighbors. By doing so, ReaGAN naturally supports both node-level personalization and autonomy (addressing Challenge 1) and joint local-global information integration (addressing Challenge 2).
2.2. Agentic Formulation
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be an attributed graph, where each node $v\in\mathcal{V}$ is associated with a text feature $t_{v}$ and an optional label $y_{v}\in\mathcal{Y}$. The goal is to predict labels $\hat{y}_{v}$ for all unlabeled nodes. Each node is treated as an agent equipped with memory $\mathcal{M}_{v}$, interacting with a frozen LLM through a multi-layered reasoning loop. At each layer, the node constructs a prompt from its previous-layer memory $\mathcal{M}_{v}^{(l-1)}$, queries the LLM for an action plan, and then updates its memory. After $L$ layers, the node queries the LLM once for prediction.
In summary, each node follows a layer-wise cycle of perception, planning, action execution, and memory update, independently deciding whether to gather local/global information, make a prediction, or take no action, and the label is predicted only at the final layer. This fully agentic workflow is detailed in Algorithm 1.
@@ 2.2. Agentic Formulation
16:   else if $a=\texttt{NoOp}$ then
17:    $\mathcal{M}_{v}^{(l)}\leftarrow\mathcal{M}_{v}^{(l-1)}$ {No change to memory}
18:   end if
19:  end for
20:end for
21:return Predicted label $\hat{y}_{v}$ (if generated)
2.3. Planning: Prompting Frozen LLM
In ReaGAN, each node is equipped with a local planner that determines how to act at each layer. Rather than relying on a globally shared aggregation rule, we leverage a frozen LLM to enable in-context planning, where decisions are conditioned on the node’s own memory. At each layer $l$, node $v$ constructs a structured prompt based on its internal memory $\mathcal{M}_{v}^{(l)}$. The prompt is then sent to a frozen LLM (e.g., LLaMA (Touvron et al., 2023), Qwen (Zhou et al., 2024), or DeepSeek (Team, 2024)), which returns an action plan:
$a_{v}^{(l)}=\texttt{LLM}(\texttt{Prompt}_{\texttt{planning}}(\mathcal{M}_{v}^{(l-1)}))$
where the action $a_{v}^{(l)}$ is selected from a discrete space and described in detail in Section 2.4. The node then parses and executes the plan, updating its memory with newly acquired information. This process is repeated for $L$ layers. At the final layer, the node constructs a prediction-specific prompt and queries the LLM to produce a label:
@@ 2.3. Planning: Prompting Frozen LLM
which concludes the planning loop illustrated in Figure 2. To sum up, the planning process allows each node to reason independently and asynchronously, determining what to do and when to act—without any global synchronization. It forms the core mechanism that enables Challenge 1 (Node-Level Autonomy) and Challenge 2 (Global Semantic Access) to be addressed in a unified and decentralized manner.
Figure 2. Overview of ReaGAN. Each node in ReaGAN is modeled as an agent equipped with four core modules: Memory, Planning, Tools, and Action. The node stores its original information and receives local and global information into its memory, constructs a natural language prompt, and queries a frozen LLM for an action plan. The planner outputs one action per layer (e.g., Local Aggregation, Global Aggregation), which are executed using available tools such as RAG. The resulting outputs are written back into memory and may produce a predicted label. This forms a closed loop of perception, planning, action, and memory refinement across layers.
2.4. Action: Node-Level Decision Space
@@
Both the aggregated summary and the selected labeled examples are written into the node’s memory, which is the same as the local version. Through this process, Global Aggregation provides semantic enrichment and label grounding from structure-agnostic sources. It expands each node’s informational horizon beyond its local neighborhood—especially benefiting nodes in sparse or isolated regions. This mechanism directly addresses Challenge 2 (Global Semantic Access).
2.4.3. NoOp.
@@
This mechanism is crucial for preventing information overload, particularly in situations where the memory already contains sufficient context. It helps prevent noise accumulation and supports pacing in multi-layer reasoning. By allowing nodes to wait or opt out of message passing altogether, NoOp reinforces ReaGAN’s core principle of self-decision and resource-aware adaptation (Challenge 1).
2.5. Memory: Tracking the Internal State
@@
where $\{m_{i}^{(l)}\}$ are the new entries produced by actions such as Local Aggregation, Global Aggregation, or retrieval-based example selection at layer $l$, where the new entries consist of $\tilde{t}_{v}^{(l)}$ and $\mathcal{E}_{v}^{(l)}$.
@@
Figure 3. Information flow from memory to prompt. Each node’s memory includes its original text feature, aggregated text feature from local and global neighbors, and selected labeled neighbor shots. During planning or prediction, these components are selectively injected into a natural language prompt, providing the LLM with (i) the input node’s raw identity, (ii) context-enhanced descriptions, and (iii) label-text pairs for few-shot learning. This design enables each agent to reason over multi-scale context and take personalized actions.
2.6. Tools: Global Semantic Augmentation